OpenStudy (anonymous):

Determine without graphing whether the given quadratic function has a maximum value, or minimum f(x)=-3x^2+12x-4

OpenStudy (anonymous):

the \(-3\) in the front should give you all you need it tells you that the parabola opens down

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-3x^2%2B12x-4 that means it has a maximum, but no minimum

OpenStudy (anonymous):

Okay I got the first solution it is max, but what is the value?

OpenStudy (anonymous):

the maximum is the second coordinate of the vertex the first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{12}{2\times (-3)}=6\] the second coordinate is what you get when you replace \(x\) by \(6\)

OpenStudy (anonymous):

okay I understand that part what I don't understand is when you take 3(2)+12(2)-4

OpenStudy (anonymous):

or 6 our answer?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i have no idea where \(3(2)+12(2)-4\) came from what you need to compute for the maximum is \[-3(6^2)+12(6)-4\]

OpenStudy (anonymous):

oh okay, so I was doing it wrong, actually I was looking the wrong problem... okay now do you start figuring out the problem, please help....